Land worth $800,000 is purchased by a down payment of $120,000 and the balance in equal monthly installments for 15 years. If interest is at j12 = 4%, determine the buyer’s and seller’s equity in the land at the end of 10 years.
Answer #1

Loan=Land Price-Down Payment=800000-120000=680000

Monthly payment=Loan*(rate/12)/(1-1/(1+rate/12)^(12*t))=680000*(4%/12)/(1-1/(1+4%/12)^(12*15))=5029.877894

Loan balance after 10 years=Loan*(1+rate/12)^(12*10)-Monthly payment/(rate/12)*((1+rate/12)^(12*10)-1)=680000*(1+4%/12)^(12*10)-5029.877894/(4%/12)*((1+4%/12)^(12*10)-1)=273117.6864

Equity of Buyer=800000-273117.6864=526882.3136

Equity of Seller=273117.6864